Given: element: 3.0 times 10 7 th atoms of Radon,
Answer #1

Half- life = 3.83 days

31 days/3.83 days = 8.093 half-life

Now we know that

After 1 half-life fraction of amount left = 1/2

After 8.093 half-life fraction left = (1/2)^8.093

atoms left = 3.0*10^7*(1/2)^8.093 = 109871.613 atoms = 1.09*10^5 atoms